OpenStudy (anonymous):

fund the exact length of the curve y=e^x, from x between 0 and 3

OpenStudy (anonymous):

\[ \Large \int_0^3 \sqrt{1+e^{2x}}dx \] yes ?

OpenStudy (anonymous):

thats part of the answer but you have to make a u substitution or two and possibly use partial fractions

OpenStudy (anonymous):

I would start with a trig substitution, maybe this will lead nowhere but let e^x=tan alpha

OpenStudy (anonymous):

will that help? \[ e^x= \tan \alpha \\ dxe^x = \sec^2{\alpha}d\alpha \longrightarrow dx=\sec^2\alpha \cot\alpha d\alpha \]